The speed of a passenger train is 14mph faster than the speed of a freight train. The passenger train travels 400miles in the same time it takes the freight train to travel 330miles. Find the speed of each train?
Found 2 solutions by jorel1380, stanbon:
Answer by jorel1380(3719)

400/n+14=330/n
400n=330n+4620
70n=4620
n=66
The freight train runs 66 mph; the passenger train runs 80 mph..

Answer by stanbon(75887)

The speed of a passenger train is 14 mph faster than the speed of a freight train. The passenger train travels 400miles in the same time it takes the freight train to travel 330miles. Find the speed of each train?
------------
Freight train DATA:
distance = 330 miles ; rate = x mph ; time = 330/x hrs
---------------------------------------
Passenger train DATA:
distance = 400 miles ; rate = x+14 mph ; time = 400/(x+14) hrs
-----------------------------------------
Equation:
time = time
330/x = 400/(x+14)
400x = 330x + 14*330
----
70x = 14*330
x = (1/5)(330)
x = 66 mph (freight speed)
x+14 = 80 mph (passenger speed)
=================================
Cheers,
Stan H.
===========